### v1

# Keep it Light! Simplifying Image Clustering via Text-Free Adapters

Abstract
Many competitive clustering pipelines have a multi-modal design, leveraging large language models (LLMs) or other text encoders, and text-image pairs, which are often unavailable in real-world downstream applications. Additionally, such frameworks are generally complicated to train and require substantial computational resources, making widespread adoption challenging. In this work, we show that in deep clustering, competitive performance with more complex state-of-the-art methods can be achieved using a text-free and highly simplified training pipeline. In particular, our approach, Simple Clustering via Pre-trained models (SCP), trains only a small cluster head while leveraging pre-trained vision model feature representations and positive data pairs. Experiments on benchmark datasets—including CIFAR-10, CIFAR-20, CIFAR-100, STL-10, ImageNet-10, and ImageNet-Dogs—demonstrate that SCP achieves highly competitive performance. Furthermore, we provide a theoretical result explaining why, at least under ideal conditions, additional text-based embeddings may not be necessary to achieve strong clustering performance in vision.
1 Introduction
Powered by the expressive capabilities of neural networks, deep clustering (DC) has redefined the state of the art (SOTA) in image clustering, outperforming traditional algorithms by indisputable margins. These DC pipelines uncover subtle features that can significantly enhance the power of downstream learners. A highly non-exhaustive list of algorithms redefining their respective SOTA includes DC-powered methods for community detection [1, 2], anomaly detection [3], image segmentation [4, 5], object detection [6, 7], and medical applications [8]. However, many recent DC pipelines rely on massive models, often powered by large language models (LLMs), which may be unnecessarily compute-intensive for clustering tasks, as well as complicated training schemes. This motivates our main research question:
Is SOTA performance achievable with a simple DC pipeline?
Indeed, we show that competitive performance, SOTA and near-SOTA, can be achieved using only a small adapter modifying the features generated by a pre-trained text-free, image encoder, i.e. not requiring any additional input text from the user, LLMs, or any other variants of text encoders. The result is a lightweight and simple end-to-end DC pipeline, which is computationally cheap enough to be run on a standard L4 GPU. Hence, our framework is easy to deploy in real-world tasks, as it is computationally accessible to most practitioners and does not require text-image pairs, which may not always be readily available in practical clustering applications.
1.1 Theoretical Motivation
The intuition behind why a text-free classifier should, at least theoretically, be able to match the power of a classifier using text embeddings together with the raw pixel data is rooted in our theoretical results. If $X$ represents a random image and $Z$ is its corresponding semantic description, which itself is a compressed representation of the information only in the random image $X$, then any classification task $Y$ that depends on both the image $X$ and its textual information can ultimately be expressed as a function of the image $X$ alone. We denote the $\sigma$-algebra generated by the random variable $X$ by $\sigma(X)$.
Proposition 1.1 (Lossless Amortization Principle (LAR)).
Fix $d,D,C\in\mathbb{N}$. Let $X,Z$ be random variables respectively, taking values in $\mathbb{R}^{d}$ and in $\mathbb{R}^{D}$, both of which are defined on a common probability space $(\Omega,\mathcal{F},\mathbb{P})$, and suppose that $Z$ is $\sigma(X)$-measurable. For every $\{0,1,\dots,C-1\}$-valued random variable $Y$ on $(\Omega,\mathcal{F},\mathbb{P})$ if: there exists a Borel measurable function $f:(\mathbb{R}^{d+D},\mathcal{B}(\mathbb{R}^{d+D}))\to(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ representing $Y$ by
$\underbrace{Y=f(X,Z)}_{\text{Depends both on $X$ and $Z$}}$ (1)
then, there is a Borel map $F:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$ providing the following lossless amortized representation
$Y=\underbrace{F(X).}_{\text{Only depends on $X$}}$ (LAR)
Proof.
See Appendix B. ∎
Further, the proposition above seems to align with the Platonic Representation Hypothesis proposed in [9], which suggests that latent representations in deep networks converge despite being trained on distinct data modalities. This supports our claim that, as long as the backbone vision model provides sufficiently good latent representations, the addition of extra modality inputs may be unnecessary.
Note that we acknowledge that, generally speaking, using multimodal representations and additional information for our networks to make predictions empirically tends to yield better results. Rather than disputing this, we aim to theorize whether, in principle and under ideal conditions, clustering could be achieved using the simplest model possible, without cumbersome multimodal data collection.
Our LAR principle (Proposition 1.1) shows that any semantically-powered image classifier may be realized by a text-free classifier “$F$”. However, can this theoretical classifier be practically realized, at-least approximately? Our main theoretical result guarantees that the theoretical text-free classifiers can indeed be approximately implemented by an MLP with a $\operatorname{softmax}$ output activation. Our guarantee holds for multiclass MLP classifiers using any standard activation function such as $\operatorname{ReLU}$, $\operatorname{Swish}$, or $\operatorname{softplus}$.
Theorem 1.2 (Text-Free DC is Powerful Enough).
In the setting of Proposition 1.1, let $\rho:\mathbb{R}\to\mathbb{R}$ be an activation function with at least one point of continuous and non-zero differentiability. Then, for every $0<\varepsilon\leq 1$, there is an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$, with activation function $\rho$, satisfying
$\big{|}\underbrace{f(X,Z)}_{\text{text-dependent classes}}-\underbrace{\,\,\operatorname{softmax}\circ\hat{F}(X)}_{\text{text-free deep learner}}\big{|}<\varepsilon$ (2)
with probability at least $1-\varepsilon$.
Proof.
See Appendix B. ∎
1.2 Simplifying Contemporary Deep Clustering
Modern clustering pipelines, e.g. [10, 11] usually involve: (i) learning powerful initial latent representations through self-supervised techniques, e.g. joint-embedding approaches [12, 13], and (ii) gradually refining the representation and clustering membership by minimizing an objective function [14].
Representation Learning in Deep Clustering
The emergence of large-scale pre-trained models optimized according to a contrastive learning objective using both natural language such as CLIP [15] and purely image-based encoders like DINO [16, 17], has proven highly effective in capturing rich, general-purpose feature representations for a variety of downstream tasks. Most SOTA-achieving DC pipelines leverage these powerful representations alongside text-embedding [18, 19] (the $Z$ in Theorem 1.2) for effective latent feature alignment. However, as we will see in Section 4, these text-based representations seem to contain redundant features, given that similar performance can be achieved without them. As we have already mentioned, being able to eliminate the need for text would make downstream applications, particularly in data-scarce regimes, much more accessible to the end user.
Self-Supervised Training
Typically, modern DC pipelines are trained using self-supervised methods by optimizing contrastive losses [20]. While contrastive methods rely on both positive and negative pairs, other approaches like BYOL [21] have demonstrated that it is possible to learn effective representations without negative pairs by simply maximizing the agreement between two views of the same data. This suggests that maximizing similarities of positive pairs alone can yield high-quality representations. We will follow a similar approach in this work too, see Section 3.
From Theory to Practice
Even if the model is approximately optimal and its parameters are perfectly adjusted, it is not clear that any real-world unsupervised training procedure can achieve the theoretical expressivity guaranteed by Theorem 1.2. Hence, we would like to empirically answer whether unimodal text-free DC pipelines can rival multimodal ones. We would like to emphasize that obtaining powerful DC performance, with the additional requirement that our DC pipeline is simple, unimodal, and text-free, is highly nontrivial since even the current SOTA in deep clustering exhibits a substantial performance gap with respect to supervised image classification models. Indeed, supervised methods typically achieve accuracies well-above $90\%$ on image datasets, while unsupervised clustering approaches rarely surpassed 50% clustering accuracy on CIFAR-20, and many fall below $50\%$ [22, 23, 24, 25]. This is because most unsupervised clustering pipelines struggle with complex natural images due to high intra-class variability.
Contributions
To address the challenges discussed so far, we propose Simple Clustering via Pre-trained models (SCP). Our main contributions are as follows:
  • Simple DC Pipeline: Our method yields competitive results while maintaining a simple yet effective architecture, requiring no additional feature layers, support sets, teacher-student networks, text information, or exponential moving averages.
  • Text-Free (Unimodal) Encodings: We integrate the powerful image encoder from CLIP (and also experiment with DINO) into a self-supervised clustering framework and show that this approach achieves competitive clustering performance on benchmark datasets such as CIFAR-10 [26], CIFAR-20 [26] and STL-10 [26] to current SOTA.
  • Principled Approach: Our text-free approach is principled, based on the approximate representation result previously presented in Theorem 1.2, which guarantees that text-free embeddings are theoretically enough.
2 Related Work
In this section, we provide a broad overview of self-supervised learning research that has inspired our work, along with recent trends in image clustering using pre-trained models.
2.1 Self-Supervised Learning
Self-supervised learning learns representations from data without explicit labels. The objective is to create a representation space where positive pairs are closer together, while negative pairs are pushed farther apart [27].
SimCLR [12] uses data augmentations, such as flipping and colour jittering, to create positive and negative pairs for optimizing objectives. It also introduces a projection head that maps embeddings into a space where contrastive loss is applied. BYOL [21] shows that high-quality representations can be learned by simply maximizing agreement between two augmented views of the same input, without requiring negative pairs. Building on these advancements, SimSiam [28] eliminates the need for both negative pairs and momentum encoders by introducing a stop-gradient operation, which effectively prevents representational collapse. Inspired by these methods, we adopt similar ideas to develop a simple and effective self-supervised framework for image clustering.
2.2 Pre-trained Models in Vision
Building on advances in self-supervised learning, CLIP [15] introduced a paradigm of contrastive pre-training that aligns images with corresponding textual descriptions. This approach enables broad task generalization without task-specific fine-tuning. DINO [16], which stands for self-distillation with no labels, demonstrates a self-supervised method for optimizing a student network from a teacher network based on vision input data only.
One of the key advantages of pre-trained models like CLIP is their ability to eliminate the need for training models from scratch for downstream tasks, significantly reducing computational costs and time. Instead of training a self-supervised neural network from the ground up, pre-trained models provide high-quality feature representations out of the box, leading to faster experimentation and improved performance on a variety of tasks. The scalability of CLIP has been further validated by openCLIP [29], which extended CLIP using the larger Vision Transformer models [30]. Similarly, models such as DINO [31] and DINOv2 [17] are capable of processing visual data and mapping it to high-quality latent representations.
2.3 Image Clustering via Pre-trained Models
To address the challenges of scaling to modern image datasets, methods such as NMCE [32] and MLC [33] have integrated deep learning with manifold clustering using the minimum coding rate principle [34]. Building on this idea, CPP [35] further refines CLIP features and estimates the optimal number of clusters when unknown. TEMI [36] improves clustering by leveraging associations between image features, introducing a variant of pointwise mutual information with instance weighting. Unlike our approach, TEMI utilizes a nearest-neighbors set and an exponential moving average for parameter optimization.
SIC [19] leverages multi-modality by mapping images to a semantic space and generating pseudo-labels based on image-semantic relationships. More recently, TAC [18] utilizes the textual semantics of WordNet [37] to enhance image clustering by selecting and retrieving nouns that best distinguish the images, facilitating collaboration between text and image modalities through mutual cross-modal neighborhood distillation.
Current pre-trained approaches often rely on heavy or complex architectures to ensure consistency, motivating us to develop a simple yet effective pipeline for image clustering. Our method requires only a simple clustering head and basic data augmentations, demonstrating strong competitiveness among recent models.
3 Our Method
Recent methods such as CC [23] and CPP [35] decouple the latent space into clustering and feature spaces. TAC [18] utilizes text information, and TEMI [36] employs self-distillation networks to enhance clustering. In contrast, our approach remains simple and efficient without relying on these techniques. As shown in Fig. 1, SCP consists of only two components: a pre-trained frozen backbone for pair construction, denoted as $f(.)$, and a trainable cluster head $g(.)$.
Figure 1: A overall pipeline for SCP. During training, two augmented views $T^{a}$ and $T^{b}$ of an image are generated from the dataset and processed by a frozen feature extractor $f$ and a trainable cluster head $g$ (a five-layer MLP). The objective is to minimize the cross-entropy loss between the outputs of the cluster head $g$ for the two augmented views.
Briefly, SCP performs data augmentations and extracts features from the augmented images using pre-trained models. The cluster head then projects these features into a cluster space, where the dimension equals the number of clusters. After training, outputs in the cluster space provide the soft assignments for clustering.
3.1 Pair Construction Backbone
The success of BYOL demonstrates that we can maximize the similarities of positive pairs without negative ones. In SCP, the positive pairs consist of samples augmented from the same instance.
Given a data instance $x_{i}$, we apply two stochastic transformations $T^{a}$ and $T^{b}$, independently selected from the augmentation family $\mathcal{T}$. This produces two correlated views: $x_{i}^{a}=T^{a}(x_{i})$ and $x_{i}^{b}=T^{b}(x_{i})$.
An appropriate augmentation strategy is vital for better downstream performance. In our work, we adopt only two simple augmentations: RandomCrop and GaussianBlur. This choice aligns with the preprocessing techniques used in training pre-trained models, ensuring compatibility with their learned representations. RandomCrop randomly crops the image to a specified size, and GaussianBlur applies a Gaussian filter to blur the image. For each image, these two augmentations are applied independently, each with a 50% probability. We then use a pre-trained model $f(\cdot)$, such as CLIP, to extract features from the augmented images: $h_{i}^{a}=f(x_{i}^{a})$ and $h_{i}^{b}=f(x_{i}^{b}).$
3.2 Cluster Head
Following the “label as representation” concept [23], when a data sample is projected into a space whose dimensionality matches the number of clusters $K$, the $k$-th component of its feature vector (after applying a $\operatorname{softmax}$ function) can be interpreted as the probability that the sample belongs to the $k$-th cluster. We employ a five-layer non-linear MLP as the clustering head $g(\cdot)$, producing a $K$-dimensional feature that is normalized with a $\operatorname{softmax}$ over the dimension of the cluster.
$y_{i}^{a}=g(h_{i}^{a}),\quad y_{i}^{b}=g(h_{i}^{b}).$
Hence, $y_{i}^{a}$ and $y_{i}^{b}$ are both $K$-dimensional vectors, whose components $y_{i,k}^{a}$ and $y_{i,k}^{b}$ indicate the probability of assigning the $i$-th sample to the $k$-th cluster. Formally, let $Y^{a},Y^{b}\in\mathbb{R}^{N\times K}$ be the outputs of the clustering head for all samples. Then, we have the following matrices:
$Y^{a}=\begin{bmatrix}y^{i}_{1}\\ \vdots\\ y^{a}_{N}\end{bmatrix}\quad Y^{b}=\begin{bmatrix}y^{b}_{1}\\ \vdots\\ y^{b}_{N}\end{bmatrix}.$
To maximize row-wise similarity, we adopt the following cross-entropy loss function instead of the commonly used InfoNCE loss [20], as SCP only have positive pairs that should share similar soft assignments:
$L_{e}=-\sum_{i=1}^{N}\sum_{k=1}^{K}y^{a}_{i,k}\log y^{b}_{i,k}.$ (3)
Inspired by the effective regularizations in TAC [18], we further introduce the following confidence loss to make the soft labels $y^{a}_{i}$ and $y^{b}_{i}$ more confident, approaching one-hot vectors:
$L_{\text{con}}=-\log\sum_{i=1}^{N}{y^{a}_{i}}^{\top}y^{b}_{i}.$ (4)
This loss ensures that the cluster head assigns higher probabilities to its top predicted clusters, thereby increasing confidence in the assignments.
In addition, following TAC [18], we introduce an entropy term $H(Y)$ to prevent model collapse, defined as follows:
$H(Y)=-\sum_{k=1}^{K}\left[P^{a}_{k}\log P^{a}_{k}+P^{b}_{k}\log P^{b}_{k}\right],$ (5)
where
$P^{a}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{a}_{i,k},\quad P^{b}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{b}_{i,k}.$
$H(Y)$ encourages uniform soft assignments across clusters, thereby mitigating the issue of empty clusters.
Hence, we define the overall objective function of SCP as
$L_{\text{clu}}=L_{e}+L_{\text{con}}-\alpha H(Y),$ (6)
where the balancing weight $\alpha$ modulates the influence of $H(Y)$, especially when the number of clusters is large. By maximizing consistency between different augmented views with regularizations, SCP effectively prevents trivial solutions and achieves competitive performance. We provide algorithm 1 to explain our pipeline.
Algorithm 1 Simple Clustering via Pre-trained Models (SCP)
1:Dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, Pre-trained model $f(\cdot)$, number of clusters $K$, batch size $B$, loss weight $\alpha$
2:Initialize cluster head $g(\cdot)$
3:for each epoch do
4:     for each mini-batch $\{x_{i}\}_{i=1}^{B}$ do
5:         Pair Construction:
6:         for each data instance $x_{i}$ in the mini-batch do
7:              Apply stochastic transformations $T^{a}$, $T^{b}$ to obtain:
8:               $x_{i}^{a}=T^{a}(x_{i})$,  $x_{i}^{b}=T^{b}(x_{i})$
9:              Extract features using pre-trained model:
10:               $h_{i}^{a}=f(x_{i}^{a})$,  $h_{i}^{b}=f(x_{i}^{b})$
11:         end for
12:         Cluster Space Encoding:
13:         for each feature $h_{i}^{a}$, $h_{i}^{b}$ do
14:              Compute soft assignments:  $y_{i}^{a}=g(h_{i}^{a})$,  $y_{i}^{b}=g(h_{i}^{b})$
15:         end for
16:         Compute Losses:
17:         Compute total clustering loss:
18:          $L_{\text{clu}}=L_{e}+L_{\text{con}}-\alpha H(Y)$
19:         Update cluster head $g(\cdot)$ parameters by minimizing $L_{\text{clu}}$
20:     end for
21:end for
22:return soft assignments $y_{i}=g(h_{i})$ for each $x_{i}\in\mathcal{X}$
4 Experiments
In this section, we evaluate the proposed SCP on six widely challenging image clustering datasets. A series of initial quantitative and qualitative comparisons, ablation studies, and hyper-parameter analyses will be carried out to investigate the effectiveness and robustness of the method.
4.1 Experimental Setup
We first introduce the datasets and metrics used for evaluation and then provide the implementation details of SCP.
4.1.1 Datasets
To evaluate the performance of SCP, we apply it to six widely used image clustering datasets: STL-10 [38], CIFAR-10 [26], CIFAR-20 [26], CIFAR-100 [26], ImageNet-10 [39], and ImageNet-Dogs [39], which is a subset from ImageNet-1k [40]. The brief information of all datasets used in our evaluation is summarized in Table 1.
Table 1: Characteristics of the benchmark datasets used in our evaluation.
Dataset Split (Train/Test) # Training # Testing # Classes
STL-10 Train/Test 5,000 8,000 10
CIFAR-10 Train/Test 50,000 10,000 10
CIFAR-20 Train/Test 50,000 10,000 20
CIFAR-100 Train/Test 50,000 10,000 100
ImageNet-10 Train/Val 13,000 500 10
ImageNet-Dogs Train/Val 19,500 750 15
4.1.2 Implementation Details
In most experiments, we utilized the CLIP [15] with the backbone ViT-B/32 [30]. For a fair comparison with TEMI [36] and CPP [35], we replaced the CLIP backbone with ViT-L/14 to match the architectures used in these methods. The cluster head $g$ is a five-layer MLP with dimensions: $D\rightarrow 1024\rightarrow 786\rightarrow 512\rightarrow 1024\rightarrow K$, where $D$ denotes the output dimension of the pre-trained model and $K$ represents the number of clusters.
We train the model using the Adam optimizer with a cosine annealing learning rate schedule, starting from an initial rate of $1\times 10^{-3}$ for 30 epochs. The batch size is set to 512 for all datasets. For regularization, we set $\alpha=2$ for CIFAR-20 and ImageNet-Dogs, $\alpha=3$ for CIFAR-100 to account for its larger number of clusters, and $\alpha=1$ for the remaining datasets. All experiments are conducted on a single NVIDIA L4 GPU. Under this setup, training SCP on CIFAR-10 takes approximately one minute, excluding data augmentation.
4.2 Main Results
We compare our method with state-of-the-art baselines on six widely-used image clustering datasets, with feature visualizations to show the competitiveness of our proposed SCP.
4.2.1 Comparison against classical and text-based methods
We first evaluate our pipeline on five widely used datasets and compare it with 18 deep clustering baselines. Since most earlier baselines adopt ResNet-34 or ResNet-18 as the backbone, we mainly focus on comparisons with CLIP and CLIP-based methods. Specifically, CLIP (zero-shot) uses CLIP’s pre-trained image and text encoders to classify images by matching them to text prompts, while CLIP (K-means) clusters directly on CLIP image features via K-means. Our approach includes SCP-CLIP, which employs a ViT-B/32 backbone like other CLIP-based methods, and SCP-DINO, which uses a ViT-B/8 backbone. Each experiment is repeated 50 times, and the best results are reported.
Table 2: Clustering performance of various baseline methods on multiple datasets (all metrics are multiplied by 100). We highlight the best and second-best results in boldfaced red and underlined in blue, respectively. CLIP (K-means), CLIP (zero-shot), SCP-CLIP, SIC, and TAC all adopt a ViT-B/32 backbone. SCP-DINO uses a Vit-B/8 backbone.
Method Text-Free STL-10 CIFAR-10 CIFAR-20 ImageNet-10 ImageNet-Dogs
NMI ACC ARI NMI ACC ARI NMI ACC ARI NMI ACC ARI NMI ACC ARI
JULE [41] 18.2 27.7 16.4 19.2 27.2 13.8 10.3 13.7 3.3 17.5 30.0 13.8 5.4 13.8 2.8
DEC [42] 27.6 35.9 18.6 25.7 30.1 16.1 13.6 18.5 5.0 28.2 38.1 20.3 12.2 19.5 7.9
DAC [39] 36.6 47.0 25.7 39.6 52.2 30.6 18.5 23.8 8.8 39.4 52.7 30.2 21.9 27.5 11.1
DCCM [43] 37.6 48.2 26.2 49.6 62.3 40.8 28.5 32.7 17.3 60.8 71.0 55.5 32.1 38.3 18.2
IIC [44] 49.6 59.6 39.7 51.3 61.7 41.1 22.5 25.7 11.7 - - - - - -
PICA [45] 61.1 71.3 53.1 59.1 69.6 51.2 31.0 33.7 17.1 80.2 87.0 76.1 35.2 35.3 20.1
CC [23] 76.4 85.0 72.6 70.5 79.0 63.7 43.1 42.9 26.6 85.9 89.3 82.2 44.5 42.9 27.4
IDFD [46] 64.3 75.6 57.5 71.1 81.5 66.3 42.6 42.5 26.4 89.8 95.4 90.1 54.6 59.1 41.3
SCAN [10] 69.8 80.9 64.6 79.7 88.3 77.2 48.6 50.7 33.3 - - - 61.2 59.3 45.7
MiCE [47] 63.5 75.2 57.5 73.7 83.5 69.8 43.6 44.0 28.0 - - - 42.3 43.9 28.6
GCC [48] 68.4 78.8 63.1 76.4 85.6 72.8 47.2 47.2 30.5 84.2 90.1 82.2 49.0 52.6 36.2
NNM [22] 66.3 76.8 59.6 73.7 83.7 69.4 48.0 45.9 30.2 - - - 60.4 58.6 44.9
TCC [24] 73.2 81.4 68.9 79.0 90.6 73.3 47.9 49.1 31.2 84.8 89.7 82.5 55.4 59.5 41.7
SPICE [25] 81.7 90.8 81.2 73.4 83.8 70.5 44.8 46.8 29.4 82.8 92.1 83.6 57.2 64.6 47.9
SIC [19] 95.3 98.1 95.9 84.7 92.6 84.4 59.3 58.3 43.9 97.0 98.2 96.1 69.0 69.7 55.8
TAC [18] 95.5 98.2 96.1 83.3 91.9 83.1 61.1 60.7 44.8 98.5 99.2 98.3 80.6 83.0 72.2
SCP-CLIP 94.6 97.9 95.1 85.5 93.2 85.7 59.5 60.1 44.3 97.4 98.5 97.9 56.3 59.6 44.6
SCP-DINO 95.8 98.4 96.5 89.2 95.2 89.8 58.9 59.3 42.8 95.9 98.8 96.7 77.1 80.5 70.5
CLIP (K-means) 92.2 94.5 89.5 71.3 75.2 62.6 50.8 48.1 30.6 96.9 98.2 96.1 39.8 38.1 20.1
CLIP (zero-shot) 93.9 97.1 93.7 80.7 90.0 79.3 55.3 58.3 39.8 95.8 97.6 94.9 73.5 72.8 58.2
(a) Step 1, $ARI=4.5$
(b) Step 10, $ARI=24$
(c) Step 50, $ARI=57$
(d) Step 100, $ARI=59$
Figure 2: Visualization of representations from different training steps learned by SCP-DINO on the ImageNet-Dogs training set, along with the corresponding clustering ARI (multiplied by 100). (a) Image embeddings directly from the DINO image encoder, with clusters obtained using K-means. (b)–(d) Image logits learned by SCP-DINO across different steps. There are $77$ steps per epoch.
As shown in Table 2, our method significantly enhances clustering performance across multiple benchmark datasets. For example, on CIFAR-10, CLIP (K-means) achieves an ACC of 75.2%, whereas SCP-CLIP improves this by 18%, surpassing TAC by 1.3%. On CIFAR-20, SCP-CLIP attains an ACC of 60.1%, outperforming SIC by 1.8%. Moreover, SCP-DINO boosts performance on four benchmarks, achieving state-of-the-art results on STL-10 and the second-best performance on both ImageNet-10 and ImageNet-Dogs. The lower performance of SCP-CLIP on ImageNet-Dogs may stem from the difficulty of capturing discriminative features in fine-grained images, particularly without aligned text features.
Notably, SCP relies exclusively on visual features but still surpasses CLIP zero-shot on all benchmarks and outperforms TAC on both STL-10 and CIFAR-10. Results underscore its effectiveness in extracting and leveraging visual information for clustering. Furthermore, our pipeline is text-free, featuring a simpler architecture with fewer hyperparameters—making it especially advantageous for adaptation to pure visual pre-trained models.
Overall, these results demonstrate that beyond zero-shot classification or multi-modal frameworks, a simpler approach can still effectively utilize the power of pre-trained models for image clustering.
4.2.2 Comparison with recent text-free methods
To supplement our answers to Question (Q1), we introduce SCP-MIX, which leverages the concatenated visual features from both CLIP and DINO. Alongside SCP-CLIP, we compare this method against two recent text-free approaches, TEMI [36] and CPP [35], as reported in Table 3.
Table 3: Clustering performance comparison for recent text-free methods. The best and second-best results are highlighted in boldface red and underlined in blue, respectively. All methods use the ViT-L/14 backbone, and SCP-MIX adopts the concatenation (or ensemble) of features learned by CLIP (ViT-L/14) and DINO (ViT-B/8).
Method Text-Free CIFAR-10 CIFAR-20 CIFAR-100
NMI ACC NMI ACC NMI ACC
TEMI [36] 92.6 96.9 64.5 61.8 79.9 73.7
CPP [35] 93.6 97.4 72.5 64.2 81.8 74.0
SCP-CLIP 93.6 97.5 69.2 65.8 80.1 74.1
SCP-MIX 95.0 98.0 72.9 67.5 80.3 73.8
SCP-MIX outperforms CPP and TEMI on both CIFAR-10 and CIFAR-20, while SCP-CLIP achieves the highest accuracy on CIFAR-100. These results underscore the competitive performance of SCP, highlighting effectiveness and simplicity. Notably, SCP achieves these outcomes without relying on additional feature layers, support sets, self-distillation, or exponential moving average strategies. Thus, unimodal text-free deep clustering method can rival multimodal pipelines.
4.3 Visualization
Figure 3: The visualization of clustering performance for SCP-CLIP with ViT-B/32 backbone. (Left): An example of an image-to-image search on STL-10, showing clusters produced by CLIP (Top) and SCP (Bottom); (Right): Visualization of clustering performance. SCP-CLIP effectively enhances CLIP’s clustering performance through a cluster head.
To provide an intuitive understanding of the clustering results, we visualize the clustering performance obtained from SCP in Fig. 3, along with learned representations on the CIFAR-10 dataset in Fig. 4. The image logits, representing SCP outputs before the final $\operatorname{softmax}$ function, are used for visualization, with t-SNE applied to reduce the feature dimensions. As shown, compared to embeddings directly learned by pre-trained models, SCP effectively forms well-separated clusters, leading to a higher ACC score. Without aligned text, SCP successfully extracts image features by incorporating a clustering head, resulting in superior within-cluster compactness and between-cluster separability. The visualization supports our quantitative results and highlights SCP’s effectiveness in learning discriminative features suitable for lightweight clustering tasks. Furthermore, Fig. 2 demonstrates that SCP requires only a few steps to establish a clustering structure, effectively separating certain image embeddings within 100 steps. For instance, from step 10 to step 50, SCP learns to push the light-green cluster away from the blue cluster.
(a) DINO, $ACC=80.5\%$
(b) SCP-Dino, $ACC=95.4\%$
(c) CLIP, $ACC=78.5\%$
(d) SCP-Clip, $ACC=93.5\%$
Figure 4: Visualization of representations learned by different methods on the CIFAR-10 training set, along with the corresponding clustering accuracy (ACC). (a) Image embeddings directly from the DINO image encoder, with clusters obtained by K-means. (b) Image logits from SCP-DINO. (c) Image embeddings directly from the CLIP image encoder, again clustered by K-means. (d) Image logits from SCP-CLIP.
4.4 Ablation Studies
To assess the effectiveness of the three loss terms, $L_{\text{e}}$, $L_{\text{con}}$, and $H(Y)$, we evaluate the performance of SCP using different combinations of these losses, as presented in Table 4. The results reveal several important insights: i) $H(Y)$ plays a crucial role in preventing cluster collapse. Without $H(Y)$, SCP tends to assign most images to only a few clusters, resulting in the model collapse on CIFAR-10 and CIFAR-20. ii) $L_{\text{con}}$ provides a slight boost in performance. The reason is that the cluster assignments would be less confident when the cluster number increases. iii) The combination of all losses effectively learns image information, leading to the best clustering performance in this table.
Table 4: Clustering results for different losses on CIFAR-10 and CIFAR-20. A ’-’ indicates model collapse.
$L_{\text{e}}$ $L_{\text{con}}$ $H(Y)$ CIFAR-10 CIFAR-20
NMI ACC ARI NMI ACC ARI
85.5 93.2 85.7 59.5 60.1 44.3
83.0 90.4 81.4 55.7 53.1 37.6
16.0 22.6 6.5 16.0 16.5 4.6
83.9 91.8 82.9 54.7 52.9 37.0
4.5 Parameter Analyses
To show how the scale of $H(Y)$ influences the performance of SCP, we evaluate it under various choices of $\alpha$ on training sets of CIFAR-10, CIFAR-20, CIFAR-100 and ImageNet-Dogs under 5 random seeds. The average results and the corresponding standard deviations are presented in Fig. 5.
Figure 5: Comparison of different loss weights $\alpha$, the solid line is the mean ARI across five runs, and the shaded region plots the standard deviation across our runs.
We observe that setting the loss weight $\alpha$ too low (e.g., below 0.5) makes it difficult for our method to converge. For CIFAR-10, the performance remains relatively stable and is not sensitive to large $\alpha$ values. In contrast, for ImageNet-dogs and CIFAR-20, choosing $\alpha=2$ leads to better results, suggesting it provides an optimal balance between regularization and scale. For CIFAR-100, which involves a larger number of clusters, a higher $\alpha$ value is more suitable. Based on these findings, we select $\alpha=3$ for CIFAR-100, $\alpha=2$ for CIFAR-20 and ImageNet-dogs, and $\alpha=1$ for all other datasets. It is also consistent with the prior knowledge of the number of clusters in each dataset.
5 Conclusion
In this work, we set out to determine whether a simple DC pipeline could achieve the performance of other recent clustering frameworks. Specifically, we aimed to achieve this without relying on text, as we argue that doing so would make image-based clustering much more practical for downstream applications. This is because image-pair datasets are highly unlikely to be readily available to practitioners for most real-world use cases. Moreover, we aimed to develop a method that is simple to train without the need for additional feature layers, support sets, teacher-student networks, or exponential moving averages.
To meet this desiderata we proposed SCP, a novel, simple, and lightweight clustering method that achieves SOTA and near-SOTA performance in several datasets such as STL-10, CIFAR-10, CIFAR-20, ImageNet-10, and ImageNet-Dogs. In particular, SCP leverages pre-trained CLIP and DINO backbones, and we anticipate these could be updated in the future with better models, or ensembles.
Limitations
Despite these promising results, several challenges remain. Firstly, our pipeline relies on data augmentations, which can be time-consuming when scaling to larger datasets such as ImageNet-1K. Also, our approach requires prior knowledge of the number of clusters, which can be difficult to determine in practical applications unless the practitioner poses strong domain knowledge. Lastly, our frameworks performance is bounded by the quality of the pre-trained model used. However, we anticipate these to continue to improve in the future.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Experiment Details
A.1 Image Search Pipeline
A.1.1 CLIP for Image search
Figure 6: CLIP Image-to-Image Search baseline depicted in Fig. 3.
Figure 6 illustrates the CLIP baseline for image-to-image search. The image repository comprises the STL-10 test split, with a target image randomly selected for the search. The search process ranks images based on Euclidean distances, displaying the top 20 most similar results. The results indicate that CLIP effectively learns discriminative features, although two mismatched images are still present in the top-20 neighborhood.
A.1.2 SCP for Image search
Figure 7 illustrates the SCP-CLIP approach for image-to-image search. SCP-CLIP successfully retrieves images from the same cluster within the top-20 neighborhood.
Figure 7: SCP-CLIP Image-to-Image Search Pipeline depicted in Fig. 3, utilizing the normalized outputs of the trained cluster head $g(.)$ without applying softmax.
A.2 Visualization
We create a t-SNE visualization image cloud of STL-10 test set to support SCP effectiveness further. It was observed that both SCP-DINO and SCP-CLIP successfully distinguished most planes, ships, horses, and deer, even when they share a blue background or similar body shape.
Figure 8: The t-SNE representations learned by SCP-DINO on STL-10. $ACC:98.4\%$
Figure 9: The t-SNE representations learned by SCP-CLIP on STL-10. $ACC:97.9\%$
Appendix B Proofs
Proof of Proposition 1.1.
By the Doob-Dyknin lemma, see e.g. [49, Lemma 1.14], since $Z$ is $\sigma(X)$-measurable if and only if there exists some Borel measurable function $g:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{D},\mathcal{B}(\mathbb{R}^{D}))$ such that
$Z=g(X).$ (7)
Now, define $F:(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))\to(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}))$ by composition, as sending every $x\in\mathbb{R}^{d}$ to
[…]
Then, together (7), the definition of $Y$ in (1), and that of $F$ in (8) yield
$Y=f(X,Z)=f(X,g(X))=F(X)$ (9)
which establishes our claim. ∎
Proof of Theorem 1.2.
[…]
$\mu(\mathcal{K}_{\varepsilon})\geq 1-\varepsilon$ (10)
and for which the restriction of the Borel measurable map $F$ to $\mathcal{K}_{\varepsilon}$ is continuous.
Now, fix a continuous activation function $\rho:\mathbb{R}\to\mathbb{R}$ with at least one point of continuous non-zero differentiability; i.e. satisfying [51, Assumption 1] (due to [52]). As shown in [51, Example 13], the softmax function satisfies [51, Assumption 8]; whence the general case of the non-Euclidean universal approximation theorem [51, Theorem 37 (ii)] applies; from which we deduce the existence of an MLP $\hat{F}:\mathbb{R}^{d}\to\mathbb{R}^{C}$ with $\rho$ activation function satisfying the uniform approximation guarantee
$\sup_{x\in\mathcal{K}_{\varepsilon}}\,|F(x)-\operatorname{softmax}\circ\hat{F}(x)|<\varepsilon.$ (11)
Define the failure set
[…]
Consequentially, the following computations are well founded: Combining (10) and (11) we deduce that
$\mathbb{P}\biggl{(}|F(X)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}$ $=\mu\biggl{(}|F(x)-\operatorname{softmax}\circ\hat{F}(x)|<\varepsilon\biggr{)}$ (12)
$=\mu(\mathbb{R}^{d}\setminus B)$
$\geq\mu(\mathcal{K}_{\varepsilon})$
$\geq 1-\varepsilon.$ (13)
Finally, by Proposition 1.1 and (1), we know that
$F(X)=f(X,Z).$ (14)
Whence, incorporating the identity in (14) into the left-hand side of the chain of inequalities in (12)-(13) implies that
$\mathbb{P}\biggl{(}|f(X,Z)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}=\mathbb{P}\biggl{(}|F(X)-\operatorname{softmax}\circ\hat{F}(X)|<\varepsilon\biggr{)}\geq 1-\varepsilon$ (15)
which concludes our proof. ∎